OpenStudy (anonymous):

Simplify the following expression (m^2)^5 / m^10

OpenStudy (anonymous):

1

OpenStudy (akshay_budhkar):

\[(x^a)^b = x^ {ab}\]

OpenStudy (anonymous):

You end up with \[\frac{m^{10}}{m^{10}} = 1\]

OpenStudy (anonymous):

when m can not be equal to 0

OpenStudy (akshay_budhkar):

so use that in this question to get \[(m^2)^5= m^{10}\]

OpenStudy (akshay_budhkar):

Yea provided m is not equal to zero

OpenStudy (anonymous):

its 1

OpenStudy (akshay_budhkar):

yes